OpenStudy (anonymous):

Build the fraction to an equivalent fraction with the specified denominator. State the numerator.. 3/7 ?/14

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Same type of question. Ans 6

OpenStudy (asnaseer):

use the fact that any number multiplied by 1 is always the same number. so x * 1 =x then use the fact that any fraction which has the same numerator and denominator is equal to 1, so:\[\frac{a}{a}=1\] so now if you a given a fraction such as:\[\frac{3}{7}\]and you are asked to convert it to:\[\frac{x}{14}\]then look at what you must multiply the original denominator (7) by to get the new denominator (14). in this it is 2. then simply multiply the original fraction by 2/2 (which is 1 and will therefore not change the result):\[\frac{3}{7}=\frac{3}{7}*\frac{2}{2}=\frac{6}{14}\]